OpenStudy (anonymous):

Solve. 0.5b+4=2(b+2)

OpenStudy (shaik0124):

0.5b+4=2b+4 sub -4 on bth side 0.5b+4-4=2b+4-4 0.5b=2b sub 0.5b on both sides 0=2b-0.5b 0=b(2-0.5) 0=b(1.5) b=0

OpenStudy (anonymous):

Would the answer be no solution? or identity

OpenStudy (shaik0124):

what else r the options

OpenStudy (anonymous):

wait nevermind I figured it out thank you so much